# Towards Interpretable Deep Generative Models via Causal Representation Learning

### v1

@@ 6.1.1 Identifiability assumptions on the mixing function
We now review approaches which place assumptions on the mixing function class $\mathcal{F}$ to identify the model.
Sparsity and anchor features.
@@
In the linear case, we saw how the “anchor features” assumption can eliminate rotation indeterminancies. Moran et al., (2022) extend the anchor feature assumption to the nonlinear case by introducing a masking variable $\bm{w}_{j}\in\{0,1\}^{K}$ for each feature $j\in[D]$:
$x_{ij}=f_{\theta}(\bm{w}_{j}\odot\bm{z}_{i})+\varepsilon_{i},$ (6.3)
@@
where $\odot$ denotes the element-wise product. The masking variable $\bm{W}=[\bm{w}_{1}^{T},\cdots,\bm{w}_{D}^{T}]$ is analogous to the sparse loadings matrix in the linear setting. With the anchor assumption, and additional assumptions on the factor covariance and parameters of $f$, Moran et al., (2022) prove that the latent factors are identifiable up to element-wise transformations and permutations. See also Xu et al., (2023), who obtain identifiability with an anchor feature assumption, as well as an additive assumption on $f$ along with independence of latent factors.
Subset condition.
@@
Around the same time, Kivva et al., (2021) proposed the “subset condition” for identifying discrete latent variables, and showed that it is nearly necessary for identification. The subset condition requires that $\operatorname{ch}(\bm{z}_{ij})$ is not a subset of $\operatorname{ch}(\bm{z}_{ik})$ for any $k\neq j$, which can be interpreted as a sparsity condition on $f_{\theta}$. In particular, anchor features are a special case of the subset condition. The basic idea of using sparsity for identifiability dates back to Pearl and Verma, (1992); Evans, (2016).
@@
There have also been recent works on nonlinear ICA which exploit sparsity in $f_{\theta}$ for identifiability (Zheng et al.,, 2022; Zheng and Zhang,, 2023), although ICA models do not consider dependence between the latent factors. These works assume a “structural sparsity” constraint on $f_{\theta}$, which is equivalent to the subset condition of Kivva et al., (2021). Also in the independent factors setting, Markham and Grosse-Wentrup, (2020); Markham et al., (2024) prove identifiability if a certain dependence graph over $\bm{x}_{i}$ admits a unique minimum edge clique cover, for which the anchor feature assumption is sufficient but not necessary.
Independent mechanisms.
@@
The principle of independent mechanisms (Section 4.3) is also related to sparsity-based approaches as well as classical assumptions from PCA. Gresele et al., (2021) operationalize this principle by assuming that the mixing function $f_{\theta}$ has a Jacobian with orthogonal columns. Gresele et al., (2021) prove that IMA eliminates certain classes of mixing functions from the set of indeterminacy transformations, $\mathcal{A}(\mathcal{F},\mathcal{P})$ when $K=D$; Ghosh et al., (2023) extend this result to $K\ll D$. In the linear case, the IMA corresponds to orthogonality conditions on the loadings matrix, as in PCA.
@@ 6.1.2 Identifiability assumptions on the latent factors
Auxiliary information.
A popular approach is to assume there is auxiliary information $\bm{u}_{i}$, which helps to “pin down” $p_{\bm{z}}$. For example, Khemakhem et al., (2020) assume that in addition to $\bm{x}_{i}$, they observe another variable $\bm{u}_{i}$ such that $p(\bm{z}_{i}|\bm{u}_{i})$ lies in an exponential family. This restriction identifies $p_{\bm{z}}$ up to a linear transformation, which can be improved to element-wise transformations when the exponential family has sufficient statistic dimension greater than two.
Mixture priors.
Kivva et al., (2022) eliminate the need for auxiliary data; instead, they assume $p_{\bm{z}}$ is a Gaussian mixture. When the clusters are distinct, this assumption prevent arbitrary rotations of $p_{\bm{z}}$, identifying the factors up to element-wise linear transformations (and permutations). Kivva et al., (2022) also prove identifiability of $f$ up to affine transformations, assuming $f$ is piecewise affine and injective (and $p_{\bm{z}}$ is a Gaussian mixture).
Temporal dependence.
Another approach is to exploit temporal structure in the latent variables. Hälvä et al., (2021) prove that $f$ is identifiable up to element-wise transform and permutation, assuming the latent distributions are stationary. More generally, Ahuja et al., (2021) consider how temporal dependencies restrict the indeterminancy set. Lachapelle et al., (2022) assume sparsity in the latent transition graph to prove element-wise and permutation identifiability. Lippe et al., (2022) assume access to known latent interventions at each time step to prove element-wise and permutation identifiability of $f$; Lippe et al., (2023) allow for unknown binary interventions.
Invariance.
Finally, identifiability can also be achieved by exploiting the existence of invariant structure in the latents under data augmentations or multiple data views. Von Kügelgen et al., (2021) assume that access to augmentations of the observed data $(\bm{x},\widetilde{\bm{x}})$ where a subset of the underlying factors for $\bm{x}$ and $\widetilde{\bm{x}}$ are the same (i.e., fixing the distribution of $\bm{z}$ in some way); this is extended to multiple augmentations in Eastwood et al., (2023). Yao et al., 2024b assume access to multiple views of the data; similarly to data augmentations, the latent variables can also be identified in this case.
@@ 6.2 Causal identifiability through interventional data
So far, we have considered the statistical identifiability of $(f,p_{\bm{z}})$. We now move onto causal identifiability of $\mathsf{G}$ and/or $A$, which requires richer types of data such as interventions. As discussed in Section 5.3, we generally cannot infer $\mathsf{G}$ from observational data alone. In fact, even when $\bm{x}_{i}$ is linear in $\bm{z}_{i}$, Squires et al., (2023) show that it is impossible to recover $\mathsf{G}$ without at least one intervention on each latent factor.
@@
This more interesting regime with interventions is called interventional causal representation learning. We now review identifiability results in interventional CRL in (i) the linear mixing case $\bm{x}_{i}=\bm{B}\bm{z}_{i}$; and (ii) the nonlinear mixing case $\bm{x}_{i}=f(\bm{z}_{i})$.
Linear mixing.
@@
Squires et al., (2023) consider the linear mixing case where $\mathcal{F}$ is the set of full rank linear transformations, parametrized by matrices in $\mathbb{R}^{D\times K}$:
$\bm{x}_{i}=\bm{B}\bm{z_{i}}+\bm{\varepsilon}_{i},i=1,\dots,N,$ (6.4)
with $\bm{z}_{i}$ satisfying the linear SEM (5.3) and where $\bm{\varepsilon}_{i}$ has independent components. (Note that $\bm{z}_{i}$ does not need to be Gaussian).
[…]
@@
The key idea behind the identification strategy of Squires et al., (2023) is that the source nodes of the graph $\mathsf{G}$ (i.e. the latent variables without any parents) leave observable implications in the data $\{\bm{x},\bm{x}^{(e)}\}$. Specifically, if $\bm{x}^{(e)}$ corresponds to a source node intervention, then the difference in the precision matrices of $\bm{x}$ and $\bm{x}^{(e)}$ will be rank-one. This is because if a source node is intervened upon, all the downstream dependencies will be the same as the observational data; only the source node distribution is modified. This strategy can be applied iteratively to identify the partial ordering (i.e. topological sort) of the latent causal graph. Now, the precision matrix of $\bm{x}_{i}^{(e)}$ satisfies $\Theta^{(e)}=(B^{\dagger})^{\top}\text{Cov}(z_{i}^{(e)})^{\dagger}B^{\dagger}$, where $\text{Cov}(z_{i}^{(e)})^{\dagger}$ is constrained due to the partial ordering of $\mathsf{G}$. This constraint allows for identifying both $B^{\dagger}$ and $\text{Cov}(z_{i}^{(e)})^{\dagger}$ from $\{\Theta,\Theta^{(e)}\}_{e\in\mathcal{E}}$. [^2] Here, $\dagger$ denotes the Moore-Penrose pseudoinverse.
Nonlinear mixing.
@@
Buchholz et al., (2023) generalize the results of Squires et al., (2023) to the nonlinear setting; specifically, they consider nonlinear mixing functions $f$ which are injective and differentiable. For the latent factors $\bm{z}_{i}$, Buchholz et al., (2023) also consider a linear SEM (5.3), but restrict $\bm{z}_{i}$ to be Gaussian. Buchholz et al., (2023) prove element-wise identifiability of the latent $\bm{z}_{i}$ as well as the causal graph of $\bm{z}_{i}$, again with the assumption of single-node perfect interventions on all the latent variables. The key idea is that the log-odds of $\bm{x}_{i}$ with respect to $\bm{x}_{i}^{(e)}$ is essentially $f^{-1}(\bm{x}_{i})^{T}(\text{Cov}(\bm{z}_{i}^{(e)})-\text{Cov}(\bm{z}_{i}))f^{-1}(\bm{x}_{i})$. Again, the difference between observational and source node precision matrices is rank-one; this geometric structure allows for $f$ to be identified up to linear transformations, at which point the results of Squires et al., (2023) can be applied.
@@
Zhang et al., (2023) also consider nonlinear mixing functions and a linear SEM on $\bm{z}_{i}$, but they allow for soft interventions instead of perfect interventions on $\bm{z}_{i}$. As soft interventions do not remove connections to parent variables, the causal graph can only be identified up to ancestors (more formally, the transitive closure; Tian and Pearl,, 2001). Consequently, Zhang et al., (2023) need an additional sparsity assumption on the graph for identifiability.
Beyond linear SEM.
@@
So far we have reviewed papers with linear SEMs on $\bm{z}_{i}$, although linearity is not strictly needed. Jiang and Aragam, (2023) study a general setting with no parametric assumptions on $\bm{z}_{i}$ and provide a minimal set of nonparametric assumptions under which the latent causal graph $\mathsf{G}$ can be identified up to so-called “isolated edges”. These assumptions are necessary in the sense that if any single assumption is violated, then counterexamples to identifiability can be constructed.
@@
One consequence of avoiding functional constraints is that statistical identifiability becomes challenging. To overcome this, Varici et al., (2024); von Kügelgen et al., (2023) require at least two hard interventions per latent node as well as “interventional discrepancy”, a requirement that the interventional distributions differ almost everywhere. von Kügelgen et al., (2023) require knowing which two interventions operate on the same latent variable; Varici et al., (2024) do not need to know which interventions are paired. The results of Varici et al., (2024); von Kügelgen et al., (2023) rely on the difference in (Stein) score functions of observational and interventional data; for estimated $\widehat{\bm{z}}=h(\bm{x})$, then $\nabla_{\widehat{z}}\log p(\widehat{z})-\nabla_{\widehat{z}}\log p^{(e)}(\widehat{z})=\bm{J}_{h^{-1}}(\widehat{z})[\nabla_{x}\log p(x)-\nabla_{x}\log p^{(e)}(x)]$. With two interventions per latent, Varici et al., (2024) learn $h$ by comparing same-node intervention score differences. Having learned $h$, the latent $\bm{z}_{i}$ and their graph can then be identified.
@@
A different line of work considers an “independent support” condition to identify latent variables (Wang and Jordan,, 2024; Ahuja et al.,, 2023). These papers make an overlap assumption in the latent space; that is, there is non-zero probability on all combinations of latent factors. To provide intuition for this assumption, it is helpful to take an experimental design perspective; we need a design table where all cells (combinations of latent factor levels) have observations. The overlap assumption means that the support of the latent factors is independent, or rectangular. Then, learning a $h$ which has independent support on $\bm{z}$ provably element-wise identifies $\bm{z}$.
7 Practical considerations
@@
In this section, we review current methodological approaches to provide an overview of how CRL can be used in practice.
@@
Variational autoencoders.  A popular approach to fit latent nonlinear representations is the variational autoencoder (Kingma and Welling,, 2014; Rezende et al.,, 2014). Consider the deep generative model (DGM):
$\bm{z}_{i}$ $\sim N(0,I)$
$\bm{x}_{i}$ $=f_{\theta}(\bm{z}_{i})+\varepsilon_{i},\quad\varepsilon_{i}\sim N(0,I).$ (7.1)
A variational autoencoder approximates the latent posterior distribution $p(\bm{z}_{i}|\bm{x}_{i})$ with the variational density:
$q(\bm{z}_{i}|\bm{x}_{i})\sim N(\mu_{\phi}(\bm{x}_{i}),\sigma_{\phi}^{2}(\bm{x}_{i})),$
[…]
@@
$\bm{z}_{i}^{(l)}=\mu_{\phi}(\bm{x}_{i})+\sigma_{\phi}(\bm{x}_{i})\odot\varepsilon_{i}^{(l)},\quad\varepsilon_{i}^{(l)}\sim N(0,I),\quad l=1,\dots,L.$
@@
Identifiable representation learning.  The DGM in (7.1) is not identifiable. As discussed in Section 6.1.1, Moran et al., (2022) introduce a masking variable $\bm{W}$ (6.3); when an anchor assumption holds, the model is identifiable. In practice, a sparsity-inducing prior $p(\bm{W})$ can be placed on the masks. The resultant ELBO is:
@@
When the prior $p(\bm{W})$ is hierarchical (e.g. the Spike-and-Slab Lasso of Ročková and George,, 2018), the additional prior parameters may be estimated using the EM algorithm (Moran et al.,, 2022). This sparse VAE can be viewed as a nonlinear analogue to factor analysis; for each column of $\widehat{W}$, the non-zero features be interpreted as a cluster.
Interventional causal representation learning.  Zhang et al., (2023) propose a method for causal representation learning from interventional data, the CausalDiscrepancy VAE. The latent $\bm{z}_{i}$ follow a linear SEM, and the observed data is a nonlinear function of the latent variables:
$\bm{z}_{i}$ $=(1-A)^{-1}\bm{\nu}_{i},\quad\bm{\nu}_{i}\sim N(0,I)$ (7.2)
$\bm{x}_{i}$ $=f_{\theta}(\bm{z}_{i})+\bm{\varepsilon}_{i},\quad\bm{\varepsilon}_{i}\sim N(0,I),$ (7.3)
where $A\in\mathbb{R}^{K\times K}$ captures the (unknown) causal graph among the $\bm{z}_{i}$. The CausalDiscrepancy VAE estimates the exogenous noise $\bm{\nu}_{i}$ distribution with the variational approximation $q(\bm{\nu}_{i}|\bm{x}_{i})\sim N(\mu_{\phi}(\bm{x}_{i}),\sigma_{\phi}^{2}(\bm{x}_{i}))$, from which $\bm{z}_{i}$ is obtained as $\widehat{\bm{z}}_{i}=(1-\widehat{A})^{-1}\widehat{\bm{\nu}}_{i}$.
[…]
@@
$\quad-\mathrm{MMD}(P_{\widehat{\theta},\widehat{\varphi}}(\widehat{X}^{(e)})||X^{(e)})-\lambda\lVert A\rVert_{1}.$ (7.5)
Zhang et al., (2023) apply the CausalDiscrepancy VAE to Peturb-Seq data (Norman et al.,, 2019). In addition to an observational dataset, there are interventional datasets, each corresponding to a single-gene CRISPR activation. The function $T_{\varphi}$ maps the intervention label (gene label) to a latent dimension; the latent dimensions can then be interpreted by inspecting which gene interventions are mapped to the same latent variable.
Score-based methods.  In Section 6.2, we discussed papers which use score functions of data to identify latent causal graphs. Two popular approaches for estimating score functions are (i) using Stein’s identity (Stein,, 1972; Rolland et al.,, 2022), and (ii) sliced score matching with variance reduction (Song et al.,, 2020).
Given the estimated data score functions, Varici et al., (2024) note that
$\nabla_{\widehat{z}}\log p(\widehat{z})-\nabla_{\widehat{z}}\log p^{(e)}(\widehat{z})=\bm{J}_{h^{-1}}(\widehat{z})[\nabla_{x}\log p(x)-\nabla_{x}\log p^{(e)}(x)],$ (7.6)
for any function $h$ with $\widehat{z}=h(X)$. These differences between the scores of the interventional and observational data should be small because the interventions act on a single-node at a time. Consequently, Varici et al., (2024) learn an $h$ that minimizes the differences in score functions. This $h$ is then used to obtain the latent estimates; the latent causal graph can also be obtained from the score function differences.

### v2

@@ 6.1.1 Identifiability assumptions on the mixing function
We now discuss approaches which place assumptions on the mixing function class $\mathcal{F}$ to identify the model.
@@
Anchor features.  In the linear case, we saw how anchor features can eliminate rotation indeterminancies. Moran et al. (2022) extend this assumption to the nonlinear case by introducing a masking variable $\bm{w}_{j}\in\{0,1\}^{K}$ for each feature $j\in[D]$:
$x_{ij}=f(\bm{w}_{j}\odot\bm{z}_{i})+\varepsilon_{ij},$ (6.3)
@@
where $\odot$ denotes the element-wise product. The masking variable $\bm{W}=[\bm{w}_{1}^{T},\cdots,\bm{w}_{D}^{T}]$ is analogous to the sparse loadings matrix in the linear setting. With the anchor assumption, and additional assumptions on the factor covariance and $f$, Moran et al. (2022) prove that the latent factors are identifiable up to element-wise transformations and permutations. See also Xu et al. (2023), who obtain identifiability with an anchor feature assumption, additionally assuming $f$ is a generalized additive model and independent latent variables. Another line of work uses anchor features (equivalent to so-called pure children) to identify latent hierarchical models via rank constraints (Huang et al., 2022; Kong et al., 2023).
@@
Subset condition.  Around the same time, Kivva et al. (2021) proposed the “subset condition” for identifying discrete latent variables in general nonlinear models, and showed that it is nearly necessary for identification. The subset condition requires that the observed children of $z_{ij}$ are not a subset of the observed children of $z_{ik}$ for any $k\neq j$, which can be interpreted as a sparsity condition on $f$. In particular, anchor features are a special case of the subset condition. This was later extended to continuous features in Kivva et al. (2022), which established the identifiability of a universal class of functions (specifically, piecewise linear $f$), addressing an open problem from Wang et al. (2021).
@@
There have also been recent works on nonlinear ICA which exploit sparsity in $f$ for identifiability (Zheng et al., 2022; Zheng and Zhang, 2023), although ICA models do not consider dependence between the latent factors. These works assume a “structural sparsity” constraint on $f$, which is equivalent to the subset condition of Kivva et al. (2021). Also in the independent factors setting, Markham and Grosse-Wentrup (2020); Markham et al. (2024) prove identifiability if a certain dependence graph over $\bm{x}_{i}$ admits a unique minimum edge clique cover, for which the anchor feature assumption is sufficient but not necessary.
@@
Independent mechanisms.  The principle of independent mechanisms (Section˜4.3) is also related to sparsity and classical assumptions from PCA. Gresele et al. (2021) operationalize this principle by assuming that the mixing function $f$ has a Jacobian with orthogonal columns, proving that this assumption eliminates certain classes of mixing functions from the set of indeterminacy transformations, $\mathcal{A}(\mathcal{F},\mathcal{P})$ when $K=D$; Ghosh et al. (2023) extend this result to $K\ll D$. In the linear case, this assumption corresponds to orthogonality conditions on the loadings matrix, as in PCA.
@@ 6.1.2 Identifiability assumptions on the latent factors
Temporal dependence.  Hyvärinen and Morioka (2016) popularized using temporal structure to identify latent factors; specifically, they assume non-stationarity of the latent factors over time to obtain identifiability for nonlinear ICA. For stationary latent distributions, Hälvä et al. (2021) prove that $f$ is identifiable up to element-wise transform and permutation. More generally, Ahuja et al. (2021) consider how temporal dependencies restrict the indeterminancy set. Lachapelle et al. (2022) assume sparsity in the latent transition graph to prove element-wise identifiability up to permutations. Lippe et al. (2022) assume access to known latent interventions at each time step to prove element-wise identifiability of $f$; Lippe et al. (2023) allow for unknown binary interventions.
Auxiliary information.  Temporal structure can be viewed as a type of auxiliary data. General auxiliary data approaches have been considered by Hyvarinen et al. (2019); Khemakhem et al. (2020), who assume that in addition to $\bm{x}_{i}$, another variable $\bm{u}_{i}$ is observed, and $p(\bm{z}_{i}|\bm{u}_{i})$ is an exponential family. This restriction identifies $p_{\bm{z}}$ up to a linear transformation, which can be improved to element-wise transformations when the exponential family has sufficient statistic dimension greater than two. Hyvarinen et al. (2019) use self-supervised learning, while Khemakhem et al. (2020) take a generative model approach.
Mixture priors.  In addition to addressing identifiability under universal approximation, Kivva et al. (2022) also eliminate the need for auxiliary data; instead, they show how mixture priors enable identification. When $p_{\bm{z}}$ is a Gaussian mixture and the clusters are distinct, this prevents arbitrary rotations of $p_{\bm{z}}$, identifying the factors up to element-wise linear transformations (and permutations). Kivva et al. (2022) also prove identifiability of $f$ up to affine transformations, assuming $f$ is piecewise affine and injective.
[…]
@@ 6.2 Causal identifiability through interventional data
So far, we have considered the statistical identifiability of $(f,p_{\bm{z}})$. We now move onto causal identifiability of $\mathsf{G}$ and/or $\bm{A}$ through the use of richer types of data such as interventions. As discussed in Section˜5.3, we generally cannot infer $\mathsf{G}$ from observational data alone. In fact, even when $\bm{x}_{i}$ is linear in $\bm{z}_{i}$, Squires et al. (2023) show that it is impossible to recover $\mathsf{G}$ without at least one intervention on each latent factor.
@@
This more difficult regime with interventions is called interventional causal representation learning. We now discuss identifiability results in interventional CRL in both the linear mixing case $\bm{x}_{i}=\bm{B}\bm{z}_{i}$, and the nonlinear mixing case $\bm{x}_{i}=f(\bm{z}_{i})$. Although we emphasize interventions here, identification through more general types of distribution shifts is an active area of research (Section 8).
@@
Linear mixing.  Squires et al. (2023) consider the linear mixing case where $\mathcal{F}$ is the set of full rank linear transformations, parametrized by matrices $\bm{B}\in\mathbb{R}^{D\times K}$:
$\bm{x}_{i}=\bm{B}\bm{z_{i}}+\bm{\varepsilon}_{i},\quad i=1,\dots,N,$ (6.4)
with $\bm{z}_{i}$ satisfying the linear SEM (5.2) and where $\bm{\varepsilon}_{i}$ has independent components. (Note that $\bm{z}_{i}$ does not need to be Gaussian). Squires et al. (2023) assume the observed data is $\{\bm{x}_{1\mathrel{\mathop{\ordinarycolon}}N},\bm{x}_{1\mathrel{\mathop{\ordinarycolon}}N_{e}}^{(e)}\}_{e\in\mathcal{E}}$ where $\bm{x}_{i}$ is drawn from the observational distribution and $\bm{x}_{i}^{(e)}$ is drawn from an interventional distribution, and prove that the latent causal graph $\mathsf{G}$ is identifiable when $\mathcal{E}$ are single-node perfect interventions acting on each of the latent variables (see Definition˜2).
@@
The key idea behind the identification strategy of Squires et al. (2023) is that the source nodes of the graph $\mathsf{G}$ (i.e. the latent variables without any parents) leave observable signatures in the data $\{\bm{x},\bm{x}^{(e)}\}$. Specifically, if $\bm{x}^{(e)}$ corresponds to a source node intervention, then the difference in the precision matrices of $\bm{x}$ and $\bm{x}^{(e)}$ will be rank-one. This is because if a source node is intervened upon, all the downstream dependencies will be the same as the observational data; only the source node distribution is modified. This strategy can be applied iteratively to identify the partial ordering (i.e. topological sort) of the latent causal graph. When the interventions are perfect, this partial ordering provides a constraint which identifies the causal graph of $\bm{z}_{i}$.
@@
Nonlinear mixing.  Buchholz et al. (2023) generalize the results of Squires et al. (2023) to the nonlinear setting; specifically, they consider nonlinear mixing functions $f$ which are injective and differentiable. For the latent factors $\bm{z}_{i}$, Buchholz et al. (2023) also consider a linear SEM (5.2), but restrict $\bm{z}_{i}$ to be Gaussian. Buchholz et al. (2023) prove element-wise identifiability of the latent $\bm{z}_{i}$, as well as the causal graph of $\bm{z}_{i}$, again with the assumption of single-node perfect interventions on all the latent variables. The key idea is the same: The difference between the observational and source node precision matrices is still rank-one; this geometric structure allows for $f$ to be identified up to linear transformations, at which point the results of Squires et al. (2023) can be applied.
@@
Zhang et al. (2023) also consider nonlinear mixing functions and a linear SEM on $\bm{z}_{i}$, but allow for soft interventions instead of perfect interventions on $\bm{z}_{i}$. As soft interventions do not remove connections to parent variables, the causal graph can only be identified up to ancestors (formally, the transitive closure; Tian and Pearl, 2001). Consequently, Zhang et al. (2023) need an additional sparsity assumption on the graph for identifiability.
@@
Beyond linear SEM.  So far we have assumed a linear SEM on $\bm{z}_{i}$, although this is not really necessary. Jiang and Aragam (2023) study a general setting with no parametric assumptions and no functional constraints, and provide a minimal set of nonparametric assumptions under which the latent causal graph $\mathsf{G}$ can be identified up to so-called “isolated edges”. These assumptions are minimal in the sense that if any single assumption is violated, then counterexamples to identifiability can be constructed.
@@
One consequence of avoiding functional constraints is that statistical identifiability becomes challenging. To overcome this, Varici et al. (2024); von Kügelgen et al. (2023) require at least two perfect interventions per latent node as well as “interventional discrepancy”, a requirement that the interventional distributions differ almost everywhere. von Kügelgen et al. (2023) require knowing which two interventions operate on the same latent variable; Varici et al. (2024) do not need to know which interventions are paired. The intuition is: for datasets arising from the same intervention, their corresponding latents must have the same distribution, except for the intervened latent dimension. This constrains the mapping from the data to the latent space, enabling identification of the latents. Then, the latent causal graph can be obtained by comparing interventional and observational data.
@@
A different line of work considers an “independent support” condition to identify latent variables (Wang and Jordan, 2024; Ahuja et al., 2023). These papers make an overlap assumption in the latent space; that is, there is non-zero probability on all combinations of latent factors. To provide intuition for this assumption, it is helpful to take an experimental design perspective; we need a design table where all cells (combinations of latent factor levels) have observations. The overlap assumption means that the support of the latent factors is independent, or rectangular. This again constrains the mapping from the data to the latent space, enabling element-wise identification of $\bm{z}$.
7 Estimation in causal representation learning
@@
In this section, we discuss current methodological approaches to estimation in order to provide an overview of how CRL can be implemented in practice.
@@
Variational autoencoders.  A popular approach to fit latent nonlinear representations is the variational autoencoder (VAE, Kingma and Welling, 2014; Rezende et al., 2014). Consider the deep generative model (DGM):
$\bm{z}_{i}$ $\sim N(\bm{0},\bm{I})$
$\bm{x}_{i}$ $=f_{\theta}(\bm{z}_{i})+\varepsilon_{i},\quad\varepsilon_{i}\sim N(\bm{0},\bm{I}).$ (7.1)
A VAE approximates the latent posterior distribution $p(\bm{z}_{i}|\bm{x}_{i})$ with the variational density:
$q_{\phi}(\bm{z}_{i}|\bm{x}_{i})\sim\mathcal{N}_{K}(\mu_{\phi}(\bm{x}_{i}),\sigma_{\phi}^{2}(\bm{x}_{i})),$
where $\mu_{\phi},\sigma_{\phi}^{2}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}^{D}\to\mathbb{R}^{K}$ are typically neural networks. Although the outputs of these neural networks are low-dimensional, the networks themselves can be arbitrarily complex and consist of any number of hidden neurons and layers.
@@
$\bm{z}_{i}^{(l)}=\mu_{\phi}(\bm{x}_{i})+\sigma_{\phi}(\bm{x}_{i})\odot\bm{\varepsilon}_{i}^{(l)},\quad\bm{\varepsilon}_{i}^{(l)}\sim\mathcal{N}_{K}(\bm{0},\bm{I}),\quad l=1,\dots,L.$
@@
Identifiable representation learning.  The DGM in (7.1) is not identifiable. As discussed in Section˜6.1.1, Moran et al. (2022) introduce a masking variable $\bm{W}$ (6.3); when an anchor assumption holds, the model is identifiable. In practice, a sparsity-inducing prior $p(\bm{W})$ can be placed on the masks. The resultant ELBO is:
@@
When the prior $p(\bm{W})$ is hierarchical (e.g. the Spike-and-Slab Lasso of Ročková and George, 2018), the additional prior parameters may be estimated using the EM algorithm (Moran et al., 2022). This sparse VAE can be viewed as a nonlinear analogue to factor analysis as $\widehat{\bm{W}}$ can be interpreted similarly to a loadings matrix: for each column of $\widehat{\bm{W}}$, the non-zero features can be considered a cluster, as they all depend on the same factor dimension.
Interventional causal representation learning.  Zhang et al. (2023) propose a method for interventional CRL, the CausalDiscrepancy VAE. The latent $\bm{z}_{i}$ follow a linear SEM, and the observed data is a nonlinear function of the latent variables:
$\bm{z}_{i}$ $=(I-\bm{A})^{-1}\bm{\nu}_{i},\quad\bm{\nu}_{i}\sim\mathcal{N}_{K}(0,I)$ (7.2)
$\bm{x}_{i}$ $=f_{\theta}(\bm{z}_{i})+\bm{\varepsilon}_{i},\quad\bm{\varepsilon}_{i}\sim\mathcal{N}_{D}(0,I),$ (7.3)
where $\bm{A}\in\mathbb{R}^{K\times K}$ captures the (unknown) causal graph among the $\bm{z}_{i}$. The CausalDiscrepancy VAE estimates the exogenous noise $\bm{\nu}_{i}$ distribution with the variational approximation $q(\bm{\nu}_{i}|\bm{x}_{i})\sim\mathcal{N}_{K}(\mu_{\phi}(\bm{x}_{i}),\sigma_{\phi}^{2}(\bm{x}_{i}))$, from which $\bm{z}_{i}$ is obtained as $\widehat{\bm{z}}_{i}=(1-\widehat{\bm{A}})^{-1}\widehat{\bm{\nu}}_{i}$.
The CausalDiscrepancy VAE also learns a map from the intervention label to the latent target and intervention shift: $T_{\varphi}\mathrel{\mathop{\ordinarycolon}}[|\mathcal{E}|]\to\{[K],\mathbb{R}\}$. This $T_{\varphi}$ allows “virtual” interventional data to be generated and compared to actual interventional data. The CausalDiscrepancy VAE combines the usual VAE objective with (i) a maximum mean discrepancy (MMD, Gretton et al., 2012) term that encourages data from the virtual interventional distribution $\widetilde{\bm{X}}^{(e)}\sim{P}_{\widehat{\theta},\widehat{\varphi}}^{(e)}$ to be close to the actual interventional data $\bm{X}^{(e)}$, and (ii) a sparsity penalty on the causal graph $\bm{A}$:
@@
$\quad-\mathrm{MMD}(\widetilde{\bm{X}}^{(e)}||\bm{X}^{(e)})-\lambda\lVert A\rVert_{1}.$ (7.5)
Zhang et al. (2023) apply the CausalDiscrepancy VAE to Peturb-Seq data (Norman et al., 2019). In addition to an observational dataset, there are interventional datasets, each corresponding to a single-gene CRISPR activation. The function $T_{\varphi}$ maps the intervention label (gene label) to a latent dimension; the latent dimensions can then be interpreted by inspecting which gene interventions are mapped to the same latent variable.
Score-based methods.  Varici et al. (2024) also propose a method for CRL from interventional data. Unlike the previous methods, Varici et al. (2024) do not use a VAE-based framework and instead rely on (Stein) score functions of the data. Specifically, for two interventions $e$ and $\widetilde{e}$, the difference in latent score functions is:
$\nabla_{\widehat{\bm{z}}}\log p^{(e)}(\widehat{\bm{z}})-\nabla_{\widehat{\bm{z}}}\log p^{(\widetilde{e})}(\widehat{\bm{z}})=\bm{J}_{h^{-1}}(\widehat{\bm{z}})[\nabla_{\bm{x}}\log p^{(e)}(\bm{x})-\nabla_{\bm{x}}\log p^{(\widetilde{e})}(\bm{x})],$ (7.6)
for any function $h$ with $\widehat{\bm{z}}=h(\bm{x})$. When $e$ and $\widetilde{e}$ intervene on the same latent dimension, the above difference is non-zero only for that dimension. This yields a constrained optimization problem for learning $h$, the solution of which identifies $\bm{z}$. The causal graph of $\bm{z}$ can then be learned by comparing score functions of observational and interventional data.